# Neural Machine Unranking

### v2

@@ IV-A Datasets
IV-B Evaluation metrics
IV-B1 Unlearning performance
@@
IV-B2 Unlearning efficiency
Unlearning efficiency is measured by the unlearn and relearn times.
Unlearn time
To ensure consistent measurement across different neural ranking models and unlearning methods we report the normalized unlearn epoch duration:
(normalized unlearn epoch duration) (11)
$\coloneqq\frac{\text{(avg time per unlearning epoch of $\mathcal{M}_{\text{{unlearn}}}$)}}{\text{(avg time per learning epoch of $\mathcal{M}_{\text{{train}}}$)}},$ (12)
@@ IV-B2 Unlearning efficiency
Relearn time
The relearn time measures the number of epochs required for the model to restore its pre-unlearning performance level. A longer relearn time is indicative of a more thorough unlearning process [13, 12, 22].
IV-C NIR models and unlearning baselines
We evaluate the proposed method alongside baseline approaches on multiple neural ranking models including two cutting-edge pretraining-based models, ColBERT [17] and BERT with Dot Productions (BERTdot) [26], along with two sophisticated word-embedding-based models, Duet [27], and MatchPyramid [28]. Table II lists the empirically chosen values of the tuning parameters $\alpha$ and $\beta$ used for each model based on performance.
TABLE II: Values of $(\alpha,\beta)$ used in the experiments.
MSMARCO query removal MSMARCO document removal TREC CAR query removal TREC CAR document removal
COLBERT (1, 5) (1, 5) (0.9, 0) (0.9, 0)
BERTdot (1, 5) (1, 5) (0.9, 0) (0.9, 0)
DUET (0.005, 0) (0.01, 0) (0.005, 0) (0.01, 0)
MatchPyramid (0.1, 0) (0.1, 0) (0.01, 0) (0.01, 0)
@@ IV-C NIR models and unlearning baselines
    Amnesiac [9, 10] continues training on $\mathcal{M}_{\text{{train}}}$ but with mislabeled samples in the forget set [^1] Graves et al. [9] proposed two unlearning methods: the first method is as described, while the second requires gradient storage during the training of $\mathcal{M}_{\text{{train}}}$ and is challenging to apply in neural ranking tasks. Following Foster et al. [10], we use only the first method and refer to it as ‘Amnesiac’.. To adapt this idea to NuMuR, we intentionally score several $\langle q,d\rangle$ pairs marked as ‘negative’ higher than those labelled as ‘positive’ in the forget set and then keep training $\mathcal{M}_{\text{{train}}}$ on the revised forget set and the original entangled set.
@@
We also report results for retrain, i.e., for retraining from scratch on only the retain set [19, 1, 20, 2] as this can be considered the ‘idealised’ approach (unless ‘controllable forgetting’ is sought). However, recall that as explained in Section II-A, obtaining $\mathcal{M}_{\text{{retrain}}}$ is typically prohibitively costly.
IV-D Experimental results
IV-D1 Unlearning performance comparison
@@
[…]
@@
Notably, in the absence of the $\mathcal{M}_{\text{{retrain}}}$, we can use the $\mathcal{M}_{\text{{train}}}$ as a surrogate for evaluating performance on the retain and test sets.
For a detailed examination across four tasks and four neural ranking models, see Figure 9 in Appendix Detailed unlearning performance. Pertinent to Goal b., additional details on the degrees of unlearning performance can be found in Figure  10 in Appendix Detailed unlearning performance.
Figure 4: Unlearning performance. The upper and lower bounds of each error bar denote the maximum and minimum MRR scores of the four tasks. See Figure 9 for detailed performance metrics for each task. Ours(1), Ours(2), and Ours(3) correspond to setting $\delta$ (i.e., the target for (6)) to half of the performance on the test set, the full performance on the test set, and the performance on the forget set, under $\mathcal{M}_{\text{{retrain}}}$, respectively. This figure illustrates that the proposed CoCoL method achieves controllable forgetting, maintaining varying degrees of performance on the forget set while maintaining performance close to the retrained model on the retain and test sets.
In terms of performance on the forget set, Amnesiac and NegGrad demonstrated the most substantial decreases. Aligning with Goal a., CoCoL consistently mirrored the performance metrics closest to the test set performance of $\mathcal{M}_{\text{{retrain}}}$.
Ideal unlearning performance on entangled and disjoint sets should mirror that of a retrained model. Therefore, we focus on Amnesiac, NegGrad, and CoCoL. As shown in Figure 4, NegGrad and Amnesiac consistently exhibited a decline in performance on both entangled and disjoint sets compared to the benchmark. In contrast, CoCoL, with different values of (6), maintained performance close to the benchmark across both sets, distinguishing it from NegGrad and Amnesiac, demonstrating controllable forgetting capacity.
For tet sets (unseen data), the performance of the retrained models serves as the benchmark, with higher performance indicating better inference ability. As shown in Figure 4, both Amnesiac and NegGrad exhibited significantly lower MRR scores compared to the benchmark, indicating their inability to maintain effective inference on unseen data. In contrast, CoCoL demonstrated not only controllable forgetting but also robust performance on unseen data.
IV-D2 Unlearning efficiency
The unlearning time for each experimental group is presented in Figure 5. When emulating the $\mathcal{M}_{\text{{retrain}}}$ (with (6) set to the performance of $\mathcal{M}_{\text{{retrain}}}$ on forget set), CoCoL consumed less time than Retrain in 13 out of 16 cases (four neural ranking models across four tasks).
Across the four neural ranking models evaluated, CoCoL demonstrated a shorter unlearning time than Retrain for the pretrained models ColBERT and BERTdot in most cases. Conversely, CoCoL was less efficient with the Duet and MatchPyramid models, indicating potential areas for improving the efficiency of CoCoL in unlearning time for conventional models.
Figure 5: Unlearn time (shorter is better). For advanced pretrained neural ranking models (ColBERT and BERTdot), CoCoL consumed significantly less time than Retrain when (6) was set to the performance on the forget set (Ours(3)), and took longer only in some groups for low forget set performance (Ours(1)). CoCoL was less efficient on pre-BERT models (MatchPyramid and Duet) compared to pretrained models. However, the former are less popular [29] and less effective (see Figure 9) than the latter.
The relearn time for each model across different tasks is detailed in Figure III, considering only (6) set to meet Goal a.. The relearn process involves iterative training the unlearned model, $\mathcal{M}_{\text{{unlearn}}}$, with the forget set until the performance on this set matches the level of the originally trained model, $\mathcal{M}_{\text{{train}}}$. Consequently, the number of epochs in this iterative training is used to represent the relearn time.
Among the four neural ranking models examined, two pretraining-based models (ColBERT and BERTdot) and Duet could be relearned within two epochs. Specifically, for the TREC CAR query removal task, the relearn time of CoCoL was two epochs, compared to one epoch for the Retrain method. For the other three task groups, both CoCoL and Retrain demonstrated a relearn time of one epoch. In the case of MatchPyramid, CoCoL outperformed Retrain in two tasks of the MS MARCO dataset, whereas Retrain excelled in the two TREC CAR tasks.
TABLE III: Relearn time (higher is better).
Model MS MARCO query removal MS MARCO document removal TREC CAR query removal TREC CAR document removal
CoCoL Retrain CoCoL Retrain CoCoL Retrain CoCoL Retrain
COLBERT 1 1 1 1 2 1 1 1
BERTdot 1 1 1 1 2 1 1 1
DUET 1 1 1 2 2 1 1 1
MatchPyramid 12 3 9 8 1 8 6 25
IV-E Loss component effectiveness and parameter sensitivity analysis
@@
To assess the effectiveness of each component in contrastive and consistent loss, we conducted an experiment with both query removal and document removal tasks from the MS MARCO dataset using the state-of-the-art ColBERT neural-ranking model.
@@
The removal of the entangled component also resulted in an accelerated forgetting rate. However, this removal significantly diminished the performance of the unlearned model on all retained and unseen data, particularly in the document removal task, where there was a marked decline in model performance across all forget, retain, and test sets after just two epochs.
@@
Figure 6: The impact of omitting different loss components. The ‘benchmark’ indicates the benchmark performance using all loss components; ‘without entangled component’ and ‘without consistent loss’ denote the performance curves when excluding loss components associated with the entangled set in (9) and the disjoint set in (10), respectively. This figure demonstrates that both the entangled and consistent loss components are crucial for balancing forgetting performance with retention and inference capabilities.
The second experiment examined the impact of parameters $\alpha$ and $\beta$ in contrastive loss, which influenced the forgetting speed and the balance between forgetting and retaining performance. Using the ColBERT model tested on the MS MARCO dataset, the results, as depicted in Figure 7, provide intuitive observations. As $\alpha$ was progressively reduced from 1 to 0.1 in document removal or from 1 to 0.5 in query removal, the model exhibited a quicker forgetting speed without significantly damaging the performance on the retain and test sets. However, smaller values of $\alpha$ (e.g., 0.05, 0.01) resulted in a dramatic decline in MRR scores in the forget set and failed to maintain performance on the retain and unseen sets. The sensitivity analysis of $\beta$ closely mirrored that of $\alpha$. An optimal setting of $\beta$ effectively accelerated forgetting.
(a) Query removal performance with varying $\alpha$.
(b) Document removal performance with varying $\alpha$.
(c) Query removal performance with varying $\beta$.
(d) Document removal performance with varying $\beta$.
Figure 7: The impact of parameters $\alpha$ and $\beta$ in the contrastive loss. The figure shows that appropriate settings for $\alpha$ and $\beta$ (specifically, $\alpha\geq 0.1$ or $\beta\leq 20$) enable effective forgetting while maintaining performance on the retain and test sets.
IV-F Discussion
Our proposed method, CoCoL, was able to achieve controllable forgetting of targeted information while preserving the overall performance of neural ranking models on retained and unseen data. However, some limitations remain. Firstly, in the unlearning performance evaluation, CoCoL exhibited shortcomings in some TREC CAR groups, indicating an inconsistency in its effectiveness across different datasets and tasks. Secondly, the necessity of manual setting for parameters $\alpha$ and $\beta$ poses a challenge. While these parameters influence the balance between forgetting and retaining performance, we provide experiential guidelines to streamline their tuning process. Finally, CoCoL does not differentiate between document removal and query removal, treating them as equivalent. A more tailored approach that recognizes and accommodates these differences could enhance the method’s precision and effectiveness. Addressing these limitations in future iterations of the method is crucial to improve its robustness and adaptability.
The exploration of CoCoL capabilities has opened avenues for future research in NuMuR. One critical area of focus should be on model-specific approaches, especially for models where CoCoL underperformed, such as MatchPyramid. CoCoL demonstrates proficiency with pretraining-based neural ranking models but the underperformance in other models suggests a need for strategies tailored to the unique characteristics of each model. Understanding and leveraging the specific features and mechanisms of different models can lead to more effective unranking approaches.
@@ IV-F Discussion
Additionally, future research should aim to distinguish between query removal and document removal more precisely. Recognizing and addressing the subtle differences between these two sub-tasks could lead to the development of more nuanced and targeted unranking methods, enhancing the overall effectiveness and accuracy of the NuMuR task.
Another significant area for advancement is the automation of parameter settings, specifically for $\alpha$ and $\beta$. Automating these settings would not only streamline the unranking process but also potentially optimize the balance between forgetting and retaining performance, making the method more accessible and flexible.
@@
Lastly, considering that neural ranking models typically comprise both embedding and ranking modules, it is imperative to investigate how unranking methods interact with these components differently. Future research should delve into the distinct impacts of unranking on embedding and ranking modules, and accordingly, develop improved unranking methods that treat these modules differently. Such an approach could lead to more effective unranking techniques, further advancing NuMuR.
V Conclusion
@@
In an era where data privacy and dynamic information landscapes are paramount, this study focuses on the field of machine unlearning, specifically within the context of neural ranking models for information retrieval (IR) systems. This research introduced the concept of Neural Machine UnRanking (NuMuR), presenting a novel method (Contrastive and Consistent Loss (CoCoL)) that effectively balances the delicate trade-off between controllable forgetting specific information and maintaining the overall performance of neural ranking models. CoCoL is particularly effective with pretraining-based neural ranking models, representing an advancement in addressing the unique challenges posed by machine unlearning in IR systems.
@@
Detailed unlearning performance
Figure 9 provides detailed unlearning performance (MRR scores of various neural ranking models across different tasks.
Figure 9: Detailed unlearning performance (MRR scores). Ours(1), (2), and (3) in legends correspond to setting (6) to half of the performance on the test set, the full performance on the test set, and the performance on the forget set, generated by CoCoL, respectively.
Figure 10 illustrates the unlearning performance of various neural ranking models across different tasks. The ‘Ideal epochs’ correspond to the stopping points that satisfy Goal a.. A clear trend is observed, where forgetting scores decline while performance on both entangled and disjoint sets remains stable (from Figure 10(a) to Figure 10(l)). Even when forgetting performance significantly diverges from test performance, CoCoL consistently maintains the stability of the retain set. This observation underscores the effectiveness of CoCoL in achieving Goal b..
In certain scenarios, CoCoL even enhances the convergence of retain sets throughout the unlearning process (as shown in Figure 10(m) and Figure 10(n)). Nevertheless, there are instances where it faces challenges in striking a precise balance between forgetting and retaining performances, evident in Figure 10(o) and Figure 10(p).
(a) MS MARCO ColBERT query removal
(b) MS MARCO ColBERT document removal
(c) TREC CAR ColBERT query removal
(d) TREC CAR ColBERT document removal
(e) MS MARCO BERTdot query removal
(f) MS MARCO BERTdot document removal
(g) TREC CAR BERTdot query removal
(h) TREC CAR BERTdot document removal
(i) MS MARCO DUET query removal
(j) MS MARCO DUET document removal
(k) TREC CAR DUET query removal
(l) TREC CAR DUET document removal
(m) MS MARCO MatchPyramid query removal
(n) MS MARCO MatchPyramid document removal
(o) TREC CAR MatchPyramid query removal
(p) TREC CAR MatchPyramid document removal

### v3

@@ 4.1 Datasets
4.2 Evaluation metrics
4.2.1 Unlearning performance
@@
4.2.2 Unlearning time
To ensure consistent measurement across different neural ranking models and unlearning methods we report the normalised unlearn epoch duration:
(normalised unlearn epoch duration) (15)
$\coloneqq\frac{\text{(avg time per unlearn epoch of $\mathcal{M}_{\text{{unlearn}}}$)}}{\text{(avg time per learning epoch of $\mathcal{M}_{\text{{train}}}$)}},$ (16)
@@ 4.2.2 Unlearning time
4.3 NIR models and unlearning baselines
We evaluate the proposed CoCoL framework on four neural ranking models for IR that leverage pretrained language models. Specifically, we consider BERTcat [26], which jointly encodes concatenated query-document pairs for relevance estimation; BERTdot [26], which encodes queries and documents separately and computes relevance via dot-product similarity; ColBERT [17], which employs contextualised late interaction to balance expressiveness and efficiency; and PARADE [27], which aggregates passage-level representations for improved document reranking.
@@ 4.3 NIR models and unlearning baselines
    Catastrophic Forgetting (CF) [28]: As a straightforward strategy, CF continues training $\mathcal{M}_{\text{{train}}}$ exclusively on the retain set, relying on the natural forgetting phenomenon in neural networks [29] to erase knowledge of the forget set.
@@
    Amnesiac [9, 10] continues training on $\mathcal{M}_{\text{{train}}}$ but with mislabelled samples in the forget set [^1] Graves et al. [9] proposed two unlearning methods: the first method is as described, while the second requires gradient storage during the training of $\mathcal{M}_{\text{{train}}}$ and is challenging to apply in neural ranking tasks. Following Foster et al. [10], we use only the first method and refer to it as ‘Amnesiac’. . To adapt this idea to NuMuR, we intentionally score several $\langle q,d\rangle$ pairs marked as ‘negative’ higher than those labelled as ‘positive’ in the forget set and then keep training $\mathcal{M}_{\text{{train}}}$ on the revised forget set and the original entangled set.
  3. 3.
@@
  4. 4.
    Selective Synaptic Dampening (SSD)[10]: SSD edits model parameters directly based on the relative importance of forget and retain samples to $\mathcal{M}_{\text{{train}}}$.
  5. 5.
    Bad Teacher (BadT) [15]: This method employs a randomly initialised model as an intentionally incompetent teacher on the forget set, while retaining $\mathcal{M}_{\text{{train}}}$ as a competent teacher for the retained set. This setup discourages the student model from preserving knowledge about the forget set during distillation.
We also report results for retrain, i.e., for retraining from scratch ($\mathcal{M}_{\text{{init}}}$) on only the retain set [19, 1, 20, 2] as this can be considered the ‘idealised’ approach (unless ‘controllable forgetting’ is sought). However, recall that as explained in Section 2.1, obtaining $\mathcal{M}_{\text{{retrain}}}$ is typically prohibitively costly.
(a) 5% Removal
(b) 15% Removal
(c) 25% Removal
Figure 4: Average unlearning performance across datasets and models for varying forget set sizes. #QR and #DR denote query removal and document removal, respectively. To better visualise forgetting ability, the forget set performance is normalised as $1-|\text{MRR}^{\mathcal{M}_{\text{{unlearn}}}}_{\text{forget}}-\text{MRR}^{\mathcal{M}_{\text{{retrain}}}}_{\text{test}}|$, where $\text{MRR}_{\mathcal{M}_{\text{{retrain}}}}^{\text{test}}$ denotes the MRR of $\mathcal{M}_{\text{{retrain}}}$ on the test set. Higher values indicate closer alignment with $\mathcal{M}_{\text{{retrain}}}$ (ideal performance on the vertex). The Y-axis ranges from 0 to 1 for all sets except the test set, which ranges from 0 to 0.5.
4.4 Experimental result
4.4.1 Unlearning performance under varying forget set sizes
@@
To evaluate the impact of varying forget set sizes, we conduct three groups of experiments where 5%, 15%, and 25% of query-document pairs are designated for forgetting, corresponding to their proportion relative to the total relevant query-document pairs.
Figure 4 summarises the results. For clarity, we report the average performance of different unlearning methods across both datasets and all four neural IR models. Detailed results for each experimental setting are provided in Table 2 and Table 3 in the Appendix.
On the forget set, consistent with Goal a., CoCoL achieves performance metrics that are most aligned with those of test set performance of $\mathcal{M}_{\text{{retrain}}}$. As depicted in Figure 4, CoCoL’s performance under both query removal and document removal settings remains closest to the ideal vertex, outperforming all baselines in maintaining this alignment. Notably, even the ‘gold standard’ Retrain [28] exhibits discrepancies between its performance on the forget set and the test set, a phenomenon observed across all three forget set sizes. For the entangled and disjoint sets, ideal unlearning requires performance comparable to $\mathcal{M}_{\text{{retrain}}}$. All baseline methods except CF consistently exhibit performance degradation on these sets relative to the benchmark. In contrast, CoCoL, maintains performance close to $\mathcal{M}_{\text{{retrain}}}$ across both sets. This stability highlights its capacity for effective forgetting.
Regarding the test set (unseen data), Retrain serves as the benchmark, where higher scores indicate better generalisation. Apart from CF, all baselines suffer notable drops in MRR, with Amnesiac exhibiting the most severe degradation. In contrast, CoCoL not only achieves superior forgetting but also sustains robust generalisation, underscoring its effectiveness in preserving inference capability on unseen data.
Figure 5: Performance shifts from $\mathcal{M}_{\text{{train}}}$ to the three forgetting destinations with a forget set size of 25%.
4.4.2 Controllable forgetting
A key property of CoCoL is its ability to achieve controllable forgetting. Specifically, by adjusting the target forget set performance specified in Eq. (6), CoCoL steers the performance of original model on the forget set toward the desired target, while maintaining stable performance on the retain and test sets.
Beyond Goal a., we further align Eq. (6) with Goal b. under two additional target values:
@@ 4.4.2 Controllable forgetting
We sort these three target values by their theoretical MRR and refer to them as destination 1 (first item of Goal b.), destination 2 (Goal a.), and destination 3 (second item of Goal b.).
We conducted experiments on the 25% forget set, with results summarised in Figure 5. Although CoCoL exhibits minor fluctuations in performance on the entangled, disjoint, and test sets across different destinations, its overall performance remains stable. This observation confirms CoCoL ’s ability to achieve controllable forgetting while preserving performance on retained and unseen data.
4.4.3 Unlearning efficiency
We further assessed the unlearning efficiency of CoCoL relative to Retrain, with results summarised in Figure 6. For destination 1, where CoCoL is expected to match Retrain’s performance across all sets, it consistently achieves substantially lower unlearn time across all tasks and models. Similarly, for destination 2, which specifies a reduced target performance, CoCoL still outperforms Retrain in terms of efficiency, underscoring its adaptability to different forgetting objectives. In contrast, for destination 3, where a substantially lower forget set performance is enforced, CoCoL requires more runtime than Retrain on most tasks.
Figure 6: Comparison of unlearning time between Retrain and CoCoL across different destinations. Here, CoCoL(D1), CoCoL(D2), and CoCoL(D3) indicate CoCoL ’s unlearning time for destinations 1, 2, and 3, respectively.
4.4.4 Ablation experiment
@@
To assess the effectiveness of each component in contrastive and consistent loss, we conducted an experiment with both query removal and document removal tasks from the MS MARCO dataset using BERTcat neural-ranking model.
@@
The removal of the entangled component also resulted in an accelerated forgetting rate. However, this removal significantly diminished the performance of the unlearned model on all retained and unseen data, particularly in the query removal task, where there was a marked decline in model performance across all forget, retain, and test sets.
@@
Figure 7: The impact of omitting different loss components. The ‘Baseline’ indicates the benchmark performance using all loss components; ‘w/o Entangled’ and ‘w/o Disjoint’ denote the performance curves when excluding loss components associated with the entangled set in (13) and the disjoint set in (14), respectively. This figure demonstrates that both the entangled and consistent loss components are crucial for balancing forgetting performance with retention and inference capabilities.
4.5 Discussion
Our experiments demonstrate that CoCoL achieves superior unlearning performance across multiple forgetting objectives, maintaining alignment with $\mathcal{M}_{\text{{retrain}}}$ on the forget set while preserving generalisation on retained and unseen data. By adjusting the target forget set performance, CoCoL exhibits controllable forgetting, enabling fine-grained trade-offs between forgetting strength and model utility. Moreover, CoCoL outperforms Retrain in efficiency, requiring substantially less runtime for most objectives. Notably, achieving stricter forgetting targets incurs higher computational costs, highlighting an inherent trade-off between forgetting precision and efficiency. These findings validate CoCoL ’s capacity as a flexible and scalable unlearning solution for neural IR systems.
4.5.1 Limitations
Although CoCoL achieves competitive unlearning performance across diverse neural IR models, it exhibits certain limitations. First, for traditional architectures such as Duet and MathPyramid, CoCoL often requires substantially more training epochs to converge and struggles to maintain stable performance on the forget set, leading to inefficiencies and suboptimal trade-offs between forgetting and retention, as shown in Fig 9 in the Appendix. Second, while CoCoL performs efficiently on PLM-based models under moderate forgetting targets, its runtime increases significantly when enforcing very low performance on the forget set, indicating a need for improved optimisation strategies under stricter constraints.
4.5.2 Future work
@@
The exploration of CoCoL capabilities has opened avenues for future research in NuMuR. One critical area of focus should aim to distinguish between query removal and document removal more precisely. Recognizing and addressing the subtle differences between these two sub-tasks could lead to the development of more nuanced and targeted unranking methods, enhancing the overall effectiveness and accuracy of the NuMuR task.
@@
In addition, considering that neural ranking models typically comprise both embedding and ranking modules, it is imperative to investigate how unranking methods interact with these components differently. Future research should delve into the distinct impacts of unranking on embedding and ranking modules, and accordingly, develop improved unranking methods that treat these modules differently. Such an approach could lead to more effective unranking techniques, further advancing NuMuR.
5 Conclusion
@@
In an era where data privacy and dynamic information landscapes are paramount, this study focuses on the field of machine unlearning, specifically within the context of neural ranking models for information retrieval (IR) systems. This research introduced the concept of Neural Machine UnRanking (NuMuR), presenting a novel method (Contrastive and Consistent Loss (CoCoL)) that effectively balances the delicate trade-off between controllable forgetting specific information and maintaining the overall performance of neural ranking models. CoCoL is effective with pretraining-based neural ranking models, representing an advancement in addressing the unique challenges posed by machine unlearning in IR systems.
Acknowledgement
This work was supported by the China Postdoctoral Science Foundation under Grant No. 2025M773205, the Postdoctoral Fellowship Program of CPSF under Grant No. GZC20252319, and the National Natural Science Foundation of China under Grant No. 72374161. The authors would also like to thank the editors and anonymous reviewers for their valuable comments and suggestions, which greatly improved the quality of this manuscript.
Data availability
To access the dataset and reproduce the experiments, please refer to the paper’s GitHub repository located at github.com/JingruiHou/NuMuR.
@@ Acknowledgement
Detailed experimental result
Table 2 and Table 3 provide comprehensive visualisations of unlearning performance for CoCoL and baseline methods across two datasets (MS MARCO and TREC CAR), three forget set sizes, and four neural IR models. Detailed performance scores for all experiments are listed in Table 4 and Table 5.
Table 2: Unlearning performance across models for varying forget set sizes on MS MARCO. #QR and #DR denote query removal and document removal, respectively. Forget set performance is normalised as $1-|\text{MRR}^{\mathcal{M}_{\text{{unlearn}}}}_{\text{forget}}-\text{MRR}^{\mathcal{M}_{\text{{retrain}}}}_{\text{test}}|$, where $\text{MRR}^{\mathcal{M}_{\text{{retrain}}}}_{\text{test}}$ is the MRR of $\mathcal{M}_{\text{{retrain}}}$ on the test set. Higher values indicate closer alignment with $\mathcal{M}_{\text{{retrain}}}$ (ideal performance on the vertex). The Y-axis ranges from 0 to 1 for all sets except the test set (0 to 0.5).
Model 5% 15% 25%
BERTcat [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
BERTdot [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ColBERT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PARADE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3: Unlearning performance across models for varying forget set sizes on TREC CAR. The radar charts use the same notation as described in the caption of Table 2.
Model 5% 15% 25%
BERTcat [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
BERTdot [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ColBERT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PARADE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Additional experiment
Figure 9 illustrates the unlearning performance of two traditonal neural ranking models (Duet and MatchPyramid) across different tasks. The ‘Ideal epochs’ correspond to the stopping points that satisfy Goal a..
While CoCoL demonstrates strong overall performance, it exhibits limitations on certain traditional models. For instance, when unlearning a trained Duet model, CoCoL requires hundreds of epochs to converge, and results in inefficiency and fluctuating performance on the forget set. Similarly, for MatchPyramid, it struggles to achieve a precise balance between forgetting and retaining, leading to suboptimal trade-offs in model utility.
(a) TREC CAR query removal with Duet
(b) TREC CAR document removal with Duet
(c) TREC CAR query removal with MatchPyramid
(d) TREC CAR document removal with MatchPyramid